A rabbit and a turtle were to start at the same place. the rabbit was feeling very confident and slept for 8 hours before starting. The turtle ran at a pace of 1 mile per hour, and the rabbit ran at 5 miles per hour. The race was a distance of 10 miles. who won the race.
Found 2 solutions by checkley79, Edwin McCravy:
Answer by checkley79(3341)

D=RT
10=1T
T=10 HOURS FOR THE TURTLE.
10=5(T-8)
10=5T-40
5T=10+40
5T=50
T=50/5
T=10 HOURS FOR THE RABBIT.
LOOKS LIKE A TIE .

Answer by Edwin McCravy(20060)

A rabbit and a turtle were to start at the same place. the rabbit was feeling very confident and slept for 8 hours before starting. The turtle ran at a pace of 1 mile per hour, and the rabbit ran at 5 miles per hour. The race was a distance of 10 miles. who won the race.
The turtle will take 10 hours to get to the finish line. 
The rabbit will take 2 hours to get to the finish line.  
While the rabbit slept the turtle traveled 8 miles. So
the turtle will make the other 2 miles in 2 hours and the
rabbit will make the whole 10 miles in those 2 hours.  So
it will end up in a tie.

Edwin